An advertising agency notices that approximately 1 in 50 potential buyers of a product sees a given magazine ad, and 1 in 8 sees a corresponding ad on television. One in 100 sees both. One in 5 actually purchases the product after seeing the ad, 1 in 10 without seeing it. What is the probability that a randomly selected potential customer will purchase the product? (Round your answer to four decimal places.)
Answer #1

Case 1: After seeing the ad AND purchasing:

(1/50 + 1/8 - 1/100) X 1/5

= (0.02 + 0.125 + 0.01) X 0.2

= 0.155 X 0.2

= 0.031

Case 2: Not seeing ad AND purchasing:

(1- 0.155) X 1/10

= 0.845 X 0.1

= 0.0845

So,

The probability that a randomly selected potential customer will purchase the product = 0.031 + 0.0845 = 0.1155

So,

Answer is:

0.1155